OpenStudy (danieke):

If a line crosses the y-axis at (0, -3) and has a slope of 3, what is the equation of the line? 3x + y = 3 3x − y = 3 -3x − y = 3 3x − y = -3

OpenStudy (kidrah69):

y=mx+b

OpenStudy (kidrah69):

m=slope

OpenStudy (kidrah69):

m=3 b=-3 now just plug in the numbers :)

OpenStudy (kidrah69):

when you have that much you should be able to figure out which equation it is .

OpenStudy (danieke):

c?

OpenStudy (kidrah69):

y=mx+b \[y=3x-3\]

OpenStudy (kidrah69):

it cant be C your slop if a POSITIVE 3

OpenStudy (danieke):

b?

OpenStudy (kidrah69):

why do you think the Y is negative?

OpenStudy (kidrah69):

no

OpenStudy (kidrah69):

do the work

TheSmartOne (thesmartone):

If you want answers, then I suggest you find another site, OpenStudy is for learning how to arrive at your answer. Not for just getting answers.

OpenStudy (firejay5):

@danieke The Correct answer is B. 3x − y = 3, because if @kidrah69 is saying it has to look like y = 3x - 3 then B. will get you there based on this: http://www.wolframalpha.com/input/?i=3x+%E2%88%92+y+%3D+3

OpenStudy (kidrah69):

@Firejay5 you should let them do the work

OpenStudy (firejay5):

@kidrah69 yes I know but dragging it on and on and on and just saying "do the work" isn't helping them at all, because you're not actually helping if you don't explain how you got to your answer.